OpenStudy (anonymous):

find a1 for a geometric sequence if r=5 and a6=1000

OpenStudy (anonymous):

\[a _{6}=ar^5\]

OpenStudy (anonymous):

substitute the values to get a

OpenStudy (anonymous):

im confused? in my class im working on geometric sequences and we have to use this formula an=a1(r)n-1

OpenStudy (anonymous):

yes according to that formular substitute a6 ie \[a _{6}=ar^{6-1}=ar^5\]

OpenStudy (anonymous):

\[1000=a(5)^5\] solve for a which is a1

OpenStudy (amistre64):

id just think about inverting the sequence: \[a_1=a_6\left(\frac 1r\right)^{6-1}\] same results tho